The product of six times a number added to twelve and fifty subtracted from two times a number is zero. Find the number(s)
Answer by ewatrrr(24785)

 
Hi
Let x and y represent the numbers
(6x +12)(2y-50) = 0
(6x +12)=0 , x = -2
(2y-50) = 0 , y = 25